a 30-60-90 triangle with hypotenuse length 28 centimeters. what is the perimeter and area of the triangle?
Answer by edjones(8007)

c=28 cm
The sides are in a ratio of 1:2:sqrt(3)
14:28:14sqrt(3)
=14:28:24.2
=14+28+24.2=66.2 cm perimeter.
A=ab/2
=14^2
=196 cm^2
.
Ed